# Toward a Well-Calibrated Discrimination via Survival Outcome-Aware Contrastive Learning

### v1

@@ C.2 Sensitive Analysis
Table 7: Impact of margin ($\alpha$) on performance metrics.
METABRIC NWTCO GBSG FLCHAIN
Methods CI IBS DDC D-CAL CI IBS DDC D-CAL CI IBS DDC D-CAL CI IBS DDC D-CAL
$\alpha$=0 0.695 0.088 0.100 0.256 0.765 0.065 0.481 0.001 0.691 0.153 0.201 0.000 0.779 0.100 0.306 0.731
$\alpha$=5 0.687 0.089 0.098 0.243 0.751 0.068 0.480 0.012 0.693 0.153 0.199 0.000 0.771 0.100 0.254 0.929
$\alpha$=7 0.691 0.088 0.083 0.134 0.773 0.059 0.463 0.503 0.695 0.151 0.182 0.000 0.783 0.090 0.280 0.693
$\alpha$=10 0.707 0.084 0.093 0.528 0.756 0.069 0.573 0.061 0.694 0.154 0.224 0.000 0.779 0.100 0.270 0.963
$\alpha$=12 0.692 0.088 0.083 0.671 0.751 0.067 0.497 0.703 0.673 0.157 0.241 0.000 0.778 0.095 0.258 0.919
$\alpha$=14 0.692 0.086 0.064 0.230 0.753 0.069 0.506 0.159 0.670 0.159 0.236 0.000 0.777 0.098 0.310 0.929
@@
Table 8: Impact of balancing coefficient ($\beta$) on performance metrics.
METABRIC NWTCO GBSG FLCHAIN
Methods CI IBS DDC D-CAL CI IBS DDC D-CAL CI IBS DDC D-CAL CI IBS DDC D-CAL
$\beta$=0.01 0.673 0.089 0.148 0.254 0.747 0.069 0.493 0.043 0.655 0.149 0.150 0.015 0.746 0.063 0.493 0.353
$\beta$=0.1 0.682 0.124 0.109 0.320 0.759 0.091 0.470 0.095 0.677 0.171 0.156 0.012 0.775 0.108 0.364 0.543
$\beta$=1.0 0.706 0.086 0.093 0.524 0.773 0.078 0.222 0.428 0.703 0.138 0.180 0.075 0.816 0.096 0.338 0.465
$\beta$=10.0 0.652 0.166 0.096 0.405 0.759 0.093 0.344 0.397 0.686 0.169 0.139 0.002 0.777 0.107 0.354 0.537
$\beta$=100.0 0.687 0.078 0.127 0.492 0.743 0.071 0.463 0.352 0.694 0.170 0.363 0.057 0.743 0.071 0.551 0.057
C.3 Effect of Contrastive Learning
Figure 6: t-SNE visualization for latent representations learned with $\mathcal{L}_{\text{{NLL}}}$ only, $\mathcal{L}_{\text{{SNCE}}}$ only, and ConSurvfor the SUPPORT and FLCHAIN dataset, colored by event times (for uncensored samples).

### v2

@@ C.3 Sensitive Analysis
Table 7: Impact of margin ($\alpha$) on performance metrics.
METABRIC NWTCO
Method CI $\uparrow$ IBS $\downarrow$ DDC $\downarrow$ D-CAL $\uparrow$ CI $\uparrow$ IBS $\downarrow$ DDC $\downarrow$ D-CAL $\uparrow$
$\alpha=0$ 0.695 0.088 0.100 0.256 0.765 0.065 0.481 0.001
$\alpha=5$ 0.687 0.089 0.098 0.243 0.751 0.068 0.480 0.012
$\alpha=7$ 0.691 0.088 0.083 0.134 0.773 0.059 0.463 0.503
$\alpha=10$ 0.707 0.084 0.093 0.528 0.756 0.069 0.573 0.061
$\alpha=12$ 0.692 0.088 0.083 0.671 0.751 0.067 0.497 0.703
$\alpha=14$ 0.692 0.086 0.064 0.230 0.753 0.069 0.506 0.159
GBSG FLCHAIN
Method CI $\uparrow$ IBS $\downarrow$ DDC $\downarrow$ D-CAL $\uparrow$ CI $\uparrow$ IBS $\downarrow$ DDC $\downarrow$ D-CAL $\uparrow$
$\alpha=0$ 0.691 0.153 0.201 0.000 0.779 0.100 0.306 0.731
$\alpha=5$ 0.693 0.153 0.199 0.000 0.771 0.100 0.254 0.929
$\alpha=7$ 0.695 0.151 0.182 0.000 0.783 0.090 0.280 0.693
$\alpha=10$ 0.694 0.154 0.224 0.000 0.779 0.100 0.270 0.963
$\alpha=12$ 0.673 0.157 0.241 0.000 0.778 0.095 0.258 0.919
$\alpha=14$ 0.670 0.159 0.236 0.000 0.777 0.098 0.310 0.929
SUPPORT SEER
Method CI $\uparrow$ IBS $\downarrow$ DDC $\downarrow$ D-CAL $\uparrow$ CI $\uparrow$ IBS $\downarrow$ DDC $\downarrow$ D-CAL $\uparrow$
$\alpha=0$ 0.592 0.205 0.150 0.000 0.883 0.008 0.962 0.999
$\alpha=5$ 0.605 0.198 0.142 0.000 0.886 0.007 0.948 0.999
$\alpha=7$ 0.604 0.202 0.145 0.000 0.874 0.012 0.984 0.991
$\alpha=10$ 0.610 0.196 0.144 0.000 0.889 0.004 0.933 0.999
$\alpha=12$ 0.608 0.197 0.147 0.000 0.879 0.008 0.954 0.999
$\alpha=14$ 0.598 0.203 0.149 0.000 0.877 0.007 0.952 0.999
Table 8: Impact of margin ($\beta$) on performance metrics.
METABRIC NWTCO
Method CI $\uparrow$ IBS $\downarrow$ DDC $\downarrow$ D-CAL $\uparrow$ CI $\uparrow$ IBS $\downarrow$ DDC $\downarrow$ D-CAL $\uparrow$
$\beta=0.01$ 0.673 0.089 0.148 0.254 0.747 0.069 0.493 0.043
$\beta=0.1$ 0.682 0.124 0.109 0.320 0.759 0.091 0.470 0.095
$\beta=1.0$ 0.706 0.086 0.093 0.524 0.773 0.078 0.222 0.428
$\beta=10.0$ 0.652 0.166 0.096 0.405 0.759 0.093 0.344 0.397
$\beta=100.0$ 0.743 0.071 0.463 0.352 0.743 0.071 0.463 0.352
GBSG FLCHAIN
Method CI $\uparrow$ IBS $\downarrow$ DDC $\downarrow$ D-CAL $\uparrow$ CI $\uparrow$ IBS $\downarrow$ DDC $\downarrow$ D-CAL $\uparrow$
$\beta=0.01$ 0.655 0.149 0.150 0.015 0.746 0.063 0.493 0.353
$\beta=0.1$ 0.677 0.171 0.156 0.012 0.775 0.108 0.364 0.543
$\beta=1.0$ 0.703 0.138 0.180 0.075 0.816 0.096 0.338 0.465
$\beta=10.0$ 0.686 0.169 0.139 0.002 0.777 0.107 0.354 0.537
$\beta=100.0$ 0.694 0.170 0.363 0.057 0.743 0.071 0.551 0.057
SUPPORT SEER
Method CI $\uparrow$ IBS $\downarrow$ DDC $\downarrow$ D-CAL $\uparrow$ CI $\uparrow$ IBS $\downarrow$ DDC $\downarrow$ D-CAL $\uparrow$
$\beta=0.01$ 0.594 0.180 0.134 0.000 0.831 0.016 0.983 0.999
$\beta=0.1$ 0.603 0.185 0.139 0.000 0.842 0.020 0.982 0.999
$\beta=1.0$ 0.610 0.183 0.132 0.000 0.854 0.010 0.975 0.999
$\beta=10.0$ 0.605 0.194 0.145 0.000 0.832 0.016 0.984 0.999
$\beta=100.0$ 0.601 0.179 0.152 0.000 0.821 0.021 0.988 0.999